OpenStudy (anonymous):

If you are asked to find 834/80, how do you know the quotient is greater than 10 before you actually divide. Please show work!

OpenStudy (anonymous):

Well first when you look at the number 834, I hope you know your multiples of 10.. Cause 80 x 10 = 800. If you break up the question.. you would see 834/80 = 800/80 + 34/80 = 10 + 34/80 <--- Simplify = 10 + 17/40 <--- Which is more than 10

OpenStudy (anonymous):

800/80=10 834>800 so 834/80>800/34 so 834/80>10